A book shop owner orders some hard-back and some soft-back versions of a book costing the shop owner $20 and $12 respectively. The total order was for 300 books.
When the order arrives there were only 200 books. The shopkeeper wishes to query the order but cannot find his copy of the original order. However, his records do tell him that it was going to cost him $5080. How many of each type of the book was his original order for?

Answer by mananth(16946)

x= cost of hardbound
y= cost of softback

1 x + 1 y = 300 .............1
Total value
20 x + 12 y = 5080 .............2
Eliminate y
multiply (1)by -12
Multiply (2) by 1
-12 x -12 y = -3600
20 x + 12 y = 5080
Add the two equations
8 x = 1480
/ 8
x = 185
plug value of x in (1)
1 x + 1 y = 300
185 + y = 300
y = 300 -185
y = 115
y = 115
x= 185 cost of hardbound
y= 115 cost of softback
m.ananth@hotmail.ca